Two wistlesA and B produce notes of frequencies 660 and 596hz eesp. There is a listener at the midpoint of the line joining them. Both the whistle B and the listener start moving with speed 30m/s away from the whistle A. If the speed of sound is 330 no. Of beats hear be lsutener is

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
 For observer note of B will not change due to zero relative motion.
Observed frequency of sound produced by A  = 660 * (330 –30)/330 = 600 Hz
 
hence no. of beats = 600 – 596 = 4
 
 
Regards
Arun (askIITians forum expert)